OpenStudy (anonymous):

complete the square on x^2+x-6=0

OpenStudy (anonymous):

\( x^2+x=6\) \(x^2+x+(\frac{1}{2})^2=6+(\frac{1}{2})^2\) \(\left(x+\frac{1}{2}\right)^2=\frac{25}{4}\) \(x+\frac{1}{2}=\pm \sqrt{\frac{25}{4}} \) \(x=-\frac{1}{2}\pm \frac{5}{2} \)

OpenStudy (anonymous):

so x = -3 or 2

OpenStudy (anonymous):

Yes.